OpenStudy (anonymous):

Solve the system: 3x^2 - 6y^2 = 9 and x^2 + 5y^2 = 53

OpenStudy (asnaseer):

do you know how to solve simultaneous equations?

OpenStudy (asnaseer):

if you do, then you could first let: a = x^2 b = y^2 and solve the system: 3a - 6b = 9 a + 5b = 53 to find a and b, then find x and y from these.